OpenStudy (anonymous):

Find the distance between the points (0, 6) and (2, 3). √13 5 √73 I keep getting -31!!

random231 (random231):

use distance formula

OpenStudy (anonymous):

thats what im doing!@

random231 (random231):

sqrt 13

OpenStudy (anonymous):

how did you get 13...I keep ending up with -31

OpenStudy (anonymous):

(2-0^2)+(3-6^2)

random231 (random231):

yea so 2^2 + 3^2

OpenStudy (anonymous):

2+3-36

random231 (random231):

then sqroot whole thing

OpenStudy (anonymous):

-33+2=-31

OpenStudy (anonymous):

x-2-x1^2+y2-y1^2

OpenStudy (anonymous):

where the hell did 4 and 9 come from??

random231 (random231):

2^2 = 4 3^2 = 9

OpenStudy (anonymous):

2-0^2+3-6^2

OpenStudy (anonymous):

2-0^2 is 2

OpenStudy (bibby):

parentheses. (2- 0)^2 = 262 = 4

OpenStudy (anonymous):

i give up

OpenStudy (anonymous):

thx anyways

OpenStudy (bibby):

Drop out

random231 (random231):

http://www.purplemath.com/modules/distform.htm

OpenStudy (anonymous):

look at how my x's are set up from yours

OpenStudy (anonymous):

you had x1-x2

OpenStudy (anonymous):

and y1-y2

random231 (random231):

yeah but thats what im telling you can always swap x1 and x2.

OpenStudy (bibby):

The order doesn't matter as long as you maintain the order